OpenStudy (anonymous):

Can you help me? Simplify x^2+13x+40/x-7 ÷ x+8/x^2-2a

OpenStudy (anonymous):

are you sure theres an (a) at the end?

OpenStudy (anonymous):

oh no yeah your right, let me fix it

OpenStudy (campbell_st):

I'd start by factoring the quadratics.... then when you divide by a fraction, you take the reciprocal and multiply

OpenStudy (anonymous):

x^2+13x+40/x-7 ÷ x+8/x^2-49

OpenStudy (campbell_st):

so can you factor the quadratics...?

OpenStudy (anonymous):

-6x3 + 13x2 + 8x - 9 / x correct?

OpenStudy (campbell_st):

no its \[\frac{(x +8)(x +5)}{(x -7)} \div \frac{(x + 8)}{(x + 7)(x - 7)}\] now apply the rule for dividing by fractions, flip and multiply \[\frac{(x + 8)(x + 5)}{(x -7)} \times \frac{(x - 7)(x + 7)}{(x + 8)}\] remove the common factors and multiply

OpenStudy (anonymous):

13x/6x × 15x something like that? or am i way off????

OpenStudy (campbell_st):

ok.... let me write it this way. remember, multiplying 2 fractions, multiply the numerators and multiply the denominators \[\frac{(x + 5)(x + 8)(x - 7)(x + 7)}{(x + 8)(x - 7)}\] what common factors do you see..?

OpenStudy (anonymous):

i see (x+8) and (x-7) as common factors. right?